OpenStudy (anonymous):

a company that packaged cold cuts needs to transport the packaged meat to the supermarket it would cost 1200 to rent a truck and it will cost an additional 100 for every 50 pounds transported write an equation that represents the transport cost for every 50 pound of cold cuts transported

OpenStudy (anonymous):

Let y equal the total cost. We know that no matter what, we are charged 1200 for the use of a truck. So we start with that. Next we let x equal the number of 50 lb "packages" we are transporting. We know that for every package, we must add 100 to y. That would give us: \[y=?x+?\]

OpenStudy (anonymous):

Where the question marks represent the respective coefficients.

OpenStudy (anonymous):

Let "y" be the total cost of transportation of the packages.............(1 Cost of renting a truck: 1200 Cost of transporting one 50 pound package=100 Let the total number of cold cut packages be "x" Cost of transporting "x" number of cold cut packages = 100*x Therefore total cost=1200+100x..................................(2) But we assumed the total cost to be "y" Hence, from (1) and (2) we get, y=100(x)+1200 ....This is the required equation :)